3 times the first of three consecutive numbers is 4 more than 2 times the second.
Answer by bhayzone(8)

Answer: 6,7,8
Let 1st num = n
2nd = n+1
3rd = n+2
3 times the first of three consecutive numbers is 4 more than 2 times the second.
3n = 2(n+1) + 4
n = 6